Circular No. 8344

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bv IN NGC 6907 S. Nakano, Sumoto, Japan, reports the discovery by R. Kushida, Yat-

sugatake South Base Observatory, of an apparent supernova (mag 15.6) on numerous unfiltered CCD survey frames taken around May 24.70 UT with a 0.40-m Schmidt-Cassegrain reflector. A follow-up image by Kushida on May 25.77 showed the new object at mag 15.3. SN 2004bv is located at α = 20^{h}25^{m}06^{s}.34, δ = −24^{o}48′53′′.7 (equinox 2000.0; measured by Y. Kushida), which is 3′′.8 west and 20′′.7 south of the center of NGC 6907. Nothing is visible at this location on earlier frames of this field back to 1995 by R. Kushida, the most recent from 2003 July 24 (limiting mag 19).

SUPERNOVA 2004bs IN NGC 3323 G. Pignata and F. Patat, European Southern Observatory; and S.

Benetti and A. Harutyunyan, Istituto Nazionale di Astrofisica, on behalf of the ‘European Research and Training Network on the Physics of Type Ia Supernovae’, report that full reduction of a spectrogram (range 380–880 nm; resolution 1.1 nm) of SN 2004bs (cf. IAUC 8341), obtained at Calar Alto Observatory on May 19.87 UT using the 2.2-m reflector, shows that this to be a type-Ib supernova, ∼ 3 weeks past maximum. The spectrum is dominated by helium lines at 587.6, 667.8 and 706.5 nm; the photospheric velocities of these three features (deduced from the minima of absorption throughs) are 10800, 10200, and 10100 km/s, respectively. Other features are identified as Ca II and Fe II. The recession velocity at the supernova position (measured from the unresolved Hα line) is 5165 km/s, in good agreement with the value for NGC 3393 listed by NED (5164 km/s).

COMET C/2003 T12 (SOHO) Further to IAUC 8341, K. Battams reports his measures for an appar-

ent comet found on SOHO C3 images by J. Danaher (discovery observation below). The reductions by R. Kisala and parabolic orbital elements by B. G. Marsden (T = 2003 Oct. 21.6 TT, q = 0.519, ω = 234^{o}.7, Ω = 181^{o}.1, i = 14^{o}.5, equinox 2000.0) appear on MPEC 2004-K33. It is possible that C/2003 T12 has a short period and somewhat smaller perihelion distance, but the latter is limited by the minimum solar elongation of 5^{o}.3; the object seems to have been too faint to show on SOHO-SWAN frames.

2003 UT  α_{2000}  δ_{2000}
Oct. 9.431  13^{h}00.^{m}9  −11^{o}39′
2004 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green